OpenStudy (anonymous):

Evaluate 4x+1/x+2 divided by 4x/x+2/x/4x+1

OpenStudy (lalaly):

\[\huge{\frac{4x+1}{x+2}\div \frac{\frac{4x}{x+2}}{\frac{x}{4x+1}}}\]\[\huge{=\frac{4x+1}{x+2} \div \frac{4x(4x+1)}{x(x+2)}}\]\[\huge{=\frac{\cancel{4x+1}}{\cancel{x+2}} \times \frac{x \cancel{(x+2)}}{4x \cancel{(4x+1)}}}\]

OpenStudy (anonymous):

1 + x^(-2) + x^(-1) + 4x

OpenStudy (anonymous):

?

OpenStudy (lalaly):

whats wrong?

OpenStudy (lalaly):

@artasena dont just write an answer like that. Show the details. ur answer is wrong anyways

OpenStudy (anonymous):

i just got confused.